Please show as much work as possible

2.0n a terrestrial planet in the Trappist-1 solar system a rocket is fired from the ground, from rest, with a combined upward acceleration of a = [21.5th(m/s ) )lk. The rocket continues with this upward acceleration for th -3.58t(m/s 2.00 s, and then ceases all thrust. From this point on e rocket is in freefall. If the total time the rocket spends in the air is 13.3 (s), then what is the acceleration of gravity on this alien world? Assume no air resistance. This is a one-dimensional problem.
Answer #1

The intiial acceleration is

a = 21.5 (2)^3 - 3.58(2)= 164.84m/s^2

The velocity of the object when it thrust is zero

V_f= at = 164.84m/s^2(2) =329.6m/s

Now applying the time of flight for the remaining journey of the rocket is

T = \frac{2 V_f }{g}

11.3 = \frac{2 (329.6)}{g}

g = 58.3 m/s^2